# Electronic structure and transport in materials with flat bands: 2D materials and quasicrystals

### v1

@@ 3 Flat bands due to static defects
Electronic transport in graphene (see for example Ref. [70, 72, 73, 75, 239, 337]) and and other graphene-related 2D materials is sensitive to static defects that are, for example, frozen ripples, screened charged impurities, or local defects like vacancies or adsorbates. Adsorbates, which can be adatoms or admolecules (H, OH, CH3, …) attached to the surface of graphene, are of great interest in the context of functionalization to control the electronic properties [79, 80, 81, 82, 338, 339, 340, 341, 84, 342, 343, 85, 344, 345, 38, 346, 347, 348, 87, 349, 337, 350]. In particular, a functionalization by an adatom (or admolecule) covalently bounded to an atom of the 2D material can be a resonant scatterer for conduction states which strongly affect electronic structure and transport properties.
@@
This appearance of a resonant state by functionalization, so-called midgap states, is particularly important in the case of bipartite lattices, i.e., systems in which a sub-network $\alpha$ of atoms are bound to atoms of a sublattice $\beta$, and reciprocally. Considering a first-neighbor tight-binding (TB) Hamiltonian, when the numbers of $\alpha$ orbitals and $\beta$ orbitals are different, isolated states appeared at the energy $E=0$ (center of the band) [351, 79, 80, 81]. Such a mechanism is exact globally for the whole structure. It also allows us to understand the appearance of resonant scatterers through a random functionalization, even if the number of $\alpha$ and $\beta$ orbitals is nearly the same. Indeed let us consider a random distribution of defects (adatoms or admolecule), with a not too strong concentration $c$ of defects. Locally, typically around a defect, an unbalance occurs between the two sublattices and a confined state appears at $E=0$ around the defect. When defects are in the same sublattice, these states are coupled to each-other and form a kind of mid-gap band, leading to a peak in the DOS of width that depends on the concentration of defects. This mechanism applies to all the 2D materials presented in this review that have bipartite lattice. Note that it gives also a simple image to understand the appearance of edge states in a bipartite lattice, since an edge necessarily produces a local unbalance between the two sublattices.
@@
Theoretical studies of transport in the presence of local defects have dealt mainly either with the Bloch-Boltzmann formalism (semi-classical methods) (see for instance [352, 239, 353, 208]). In these theories, a major length scale that characterizes the electron scattering is the elastic mean-free path $L_{e}$. These approaches indeed explain some experimental observations such as the quasi-linear variation of conductivity with a concentration of charge carriers. Yet, these theories have important limitations and can hardly describe in detail the localization phenomena that have been reported in some experiments. Indeed, in the presence of a short range potential, such as that produced by local defects, the electronic states are localized on a length scale $\xi$. A sample will be insulating unless some source of scattering, like electron-electron or electron-phonon interaction, leads to a loss of the phase coherence on a length scale $L_{i}$, $L_{i}<\xi$. Therefore, in addition to the elastic meanfree path $L_{e}$, the inelastic meanfree path $L_{i}$ and the localization length $\xi$ also play a fundamental role for the conductivity of graphene with adsorbates. Moreover, the study of transport by semi-classical methods is valid when Fermi energy $E_{F}$ is far enough from the Dirac energy. But for $E_{F}$ close to Dirac energy, the effects of resonant scatterers can strongly modify the electronic structure (such as gap opening [354, 160]…). These effects are important when the resonant scatterer concentration (defect concentration) is large with respect to the charge carrier concentrations; indeed, each resonant scatterer creates one midgap states at Dirac energy $E_{D}$ [79, 80, 81]. Since these quantum effects, beyond the semi-classical behavior, are extremely dependent on the type of functionalization, systematic theoretical studies are needed to understand experimental results and stimulate new experimental studies.
@@ 3.1 Quantum transport calculation and calculated quantities
[…]
@@
At finite temperature $T$, inelastic scattering caused by electron-phonon interactions are implanted by using the approximation of Relaxation Time Approximation (RTA). The details of the implementation of the RTA are given in [38] (A.1.2). The conductivity $\sigma$ and the inelastic mean free path $L_{i}$ in the $x$-direction are thus calculated at every energy $E$ and every inelastic scattering time $\tau_{i}$. $\tau_{i}$ is the time beyond which the velocity autocorrelation function (Appendix Eq. (62)) goes exponentially to zero, and $L_{i}(E_{F},\tau_{i})$ is the typical distance of propagation during the time interval $\tau_{i}$ for electrons at energy $E$. Thus $L_{i}$ is also the distance beyond which a wave packet loses its phase coherence due to inelastic scattering processes, whereas elastic scattering events do not destroy the phase coherence. We know that $L_{i}$ decreases when the temperature $T$ increases, however the exact dependence of $L_{i}$ versus $T$ is unknown. This is why we consider different cases according to different possible values of $L_{i}$. Indeed, three different transport regimes may exist depending on $L_{i}$ value with respect to the elastic mean free path $L_{e}$, which is the average distance between two elastic scattering events. $L_{e}$ can be calculated approximately using phenomenological formula (67) of the Appendix. The different transport regimes are:
@@
    Localization regime. When $L_{i}\gg L_{e}$, multiple scattering effects (such as weak or strong localization) strongly affect the transport and the conductivity is “macroscopic” in the sense that it is established over large sample sizes. This happens at sufficiently low temperature $T$, and then $\sigma$ decreases when $L_{i}$ increases (i.e., $T$ decreases). At each energy $E$, the localization length $\xi$ can be extracted by extrapolation of the curve $\sigma$ versus $L_{i}$ curves, when $\sigma(L_{i}=\xi)=0$.
@@
    Diffusive regime. For smaller $L_{i}$ values, since $L_{i}\gtrsim L_{e}$, i.e., at higher temperatures (typically room temperature), $\sigma(L_{i})$ reaches a conductivity “plateau” close to the maximum $\sigma$ value (see A.1.2),
@@
    This regime is called “diffusive” regime, where $\sigma(L_{i})$ is almost independent on $L_{i}$ over a large $L_{i}$ range depending on the energy $E_{F}$. Usually the conductivity plateau corresponds to $L_{i}$ values from a few nm to a few 10 nm, which may correspond to high temperatures and room temperatures, respectively. In this case, the conductivity of a sample depends only on quantum scattering over small distances which are typically of the order of magnitude of the distances between static defects ($L_{e}$); this is the reason why $\sigma_{M}$ is called “microscopic” conductivity. At each energy, the microscopic diffusivity $\mathcal{D}_{M}$ and microscopic conductivity $\sigma_{M}$ are defined as the maximum value of $\mathcal{D}(\tau_{i})$ and $\sigma(\tau_{i})$, respectively. The corresponding diffusivity is also called the microscopic diffusivity, $\mathcal{D}_{M}$, and it is possible to deduce from it the elastic mean-free path $L_{e}$ (Eq. (67)).
@@
    Ballistic regime. The situation $L_{i}<L_{e}$ is an extreme case that most of time should not be reached often in real materials at realistic temperature (room temperature). This corresponds to the case of very pure materials with very few static defects. The conductivity is independent of static defects, and thus $\sigma(L_{i})$ increases when $L_{i}$ increases.
@@ 3.2 Mono and bilayer graphene with random distribution of defects
The tight binding (TB) Hamiltonian is,
@@
where $t_{ij}$ are the hopping terms between $p_{z}$ orbitals $i$ and $j$, and $\langle i,j\rangle$ is the double sum on index $i$ and $j$ with $i\neq j$. We can include only the first neighbor hopping terms (first-neighbor TB: intralayer hopping $t_{ij}=-t_{0}=-2.7$ eV and for Bernal bilayer graphene (BLG) interlayer hopping $t_{ij}=t_{1}=0.34$ eV), or hopping terms beyond the first neighbors which is often more realistic using the same SK-TB parameters as in Sec. 2.3.2. Both models correctly recover the monolayer graphene (MLG) and Bernal bilayer graphene (BLG) energy band dispersions in the vicinity of the Dirac energy level $E_{D}=0$ [94, 354]. For MLG and BLG, the spectrum is symmetric with respect to $E_{D}$ (electron-hole symmetry) for first-neighbor TB (see Sec. 3.3.1), whereas it is slightly di-symmetrical due to coupling beyond the first neighbors with the SK-TB (Sec. 2.3.2). It is important to note that another TB model exists for the BLG, in particular a well-known model deduced from TB model for the bulk graphite (see Ref. [208] and Refs therein). For the properties that we are studying here, the latter gives results comparable to the SK-TB model (Sec. 2.3.2).
@@ 3.2.1 Monolayer graphene (MLG)
Electronic properties in graphene with static defects such as example vacancies, substitutions or adsorbates has been intensively studied (see for instance Refs. [79, 82, 83, 374, 84, 345, 357, 346, 38, 358, 354, 375, 160]). We can not go into all these studies in detail here, but we will just mention a few aspects that we consider essential. Roughly speaking, it is possible to propose a unified description of transport in one graphene sheets with adsorbates that fully takes into account localization effects and loss of electronic coherence due to inelastic processes [38]. For this purpose, we have distinguished between defects that are resonant scatterers or defects that are non-resonant scatterers. Examples of resonant scatterers are adsorbates that create a covalent bond with a carbon atom of a graphene layer, such as simple atoms or molecules adsorbate: H, OH, CH3… To simulate this covalent bond, we assume that the $p_{z}$ orbital of carbon, just below the adsorbate, is removed [81, 338, 339]. As explained in the introduction of this part. 3, in a monolayer graphene (MLG), which is a zero-gap material with a bipartite lattice, such functionalization states create so-called “midgap states” at the Dirac energy $E_{D}=0$. If we imagine a periodic structure with a low concentration of resonant defects, so that the defect states do not interact with each other, the midgap states would be associated with strict flat bands. These midgap states at $E_{D}$ energy are states that are only uncoupled from the rest of the band if we consider a Hamiltonian that includes only first-neighbor couplings. If couplings beyond the first neighbors are taken into account, the midgap states are coupled into themselves, and we then obtain a “midgap peak” whose width depends on the defect concentration, at an energy shifted with respect to $E_{D}$[372, 358]. This midgap peak is associated to flat bands in a periodic system. Examples of non-resonant adsorbates is adsorbates that create two covalent bounds with two neighboring C atoms of the MLG. This is the case for the adsorption of an oxygen atom (epoxy group). In generic terms, a non-resonant disorder can also be simulated by an Anderson disorder (random variation of the on-site energy of all orbitals). Such defects do not create resonant peaks, so they do not alter strongly the DOS, but they do affect the transport properties of charge carriers [38]. As non-resonant defects are not associated with flat bands, we will not go into too much detail in this review.
@@ 3.2.2 Bernal bilayer graphene (BLG)
Figure 14: Electronic properties of Bernal AB bilayer graphene (BLG). (a) Sketch of the crystal structure of BLG. Atoms A1 and B1 are on the lower layer (layer 1); A2 , B2 on the upper layer (layer 2). First-neighbor TB: $V^{0}_{pp\pi}=-t_{0}=-2.7$ eV and $V^{0}_{pp\sigma}=t_{1}=0.34$ eV are the intralayer and interlayer first-neighbor hopping terms, respectively. BLG is bipartite lattice between sublattices $\alpha=\{{\rm A}_{1},{\rm B}_{2}\}$ and $\beta=\{{\rm A}_{2},{\rm B}_{1}\}$. Total DOS of (b) Monolayer (MLG) and (c) BLG with random distribution of vacancies with concentrations $c$ (%) of vacant atoms: (dashed line) First-neighbor TB (see panel (a)), and (solid line), SK-TB with hopping beyond first neighbor (see Sec. 2.3.2). The black dashed curve is pristine BLG ($c=0$), calculated with first-neighbor TB. (d) Conductivity $\sigma$ (solid line) and inelastic scattering length $L_{i}$ (dashed line) in BLG versus inelastic scattering time $\tau_{i}$ for concentration $c=1\%$ of resonant adsorbates (vacancies), and for 3 energy values, calculated with first-neighbor TB. (e,f) Microscopic conductivity $\sigma_{M}$ calculated with first-neighbor TB: (e) $c<c_{l}\sim 1.5$% and $\sigma_{M}$ with a constant minimum value (“plateau”) which decreases as $c$ increases, (f) $c>c_{l}\sim 1.5$% and $\sigma_{M}$ reaches a minimum value independent of $c$ value (see text). Concentration of vacancies: semi circle $c=0.05\%$, cross $c=0.25\%$, triangle up $c=0.5\%$, circle $c=1\%$, rectangle $c=2\%$, square $c=3\%$, triangle down $c=5\%$, star $c=10\%$. Insert shows enlarged curves around $E_{D}=0$. $G_{0}={2e^{2}}/{h}$. Numerical method see A.3. From [358, 376].
@@
The unit cell of Bernal (BLG) contains four carbon atoms, A1, B1 in layer 1 and A2, B2 in layer 2 (Fig. 14). Atoms A have three B first neighbors in the same layer and one A neighbor in the other layer, while atoms B have only three A first neighbors in the same layer. Since BLG is a bipartite lattice (see Sec. 16), an isolated functionalization by resonant scatterers create an midgap state like in monolayer graphene (MLG). We first consider now the case of a random distribution of resonant defects in the two layer in BLG [358]. Figure 14 shows the total density of states (total DOS) $n(E)$ for different concentrations $c=0.5\%$ to $10\%$ of defects in MLG and BLG.
@@
Without defects, $c=0$, MLG and BLG DOSs differ [38, 358] for energies $E$ such as $-t_{1}<E<t_{1}$ [208], where $t_{1}$ is the interlayer hopping term (Fig. 14(b,c)). For small $c$ concentrations, $c<c_{l}\approx 1.5\%$, this distinction is still observed. But remarkably, MLG and BLG have very similar DOS, for concentrations of defects $c$ larger than $c_{l}$. With TB model (first-neighbor hopping only), midgap states occur at energy $E_{MLG}=0$. With the TB model including all neighbors [358], the midgap state is no longer at $E=0$, but it becomes a broad peak at negative energy $E_{MLG}$. $E_{MLG}$ value varies from $E_{MLG}\approx-0.2\,eV$ to $E_{MLG}\approx-0.3\,eV$ when defect concentration increases from $c=0.5\%$ to $10\%$, as in monolayer graphene [239, 372]. For large values of energies $E$ the DOS is weakly affected by the presence of disorder and near the Dirac energy there is an intermediate regime where the pseudo-gap is filled (“plateau”). We distinguish three cases according to energy values, which correspond to different transport regimes (Fig. 14) [358]:
@@
    For energies $E$ in the DOS “plateau” due to vacancies but not in the midgap states $E\neq E_{MLG}$ (Fig. 14(a)), the microscopic conductivity ($\sim$ room temperature conductivity) $\sigma_{M}$ (Fig. 14(e,f)) shows a “plateau” independent of $E$. For small $c$, $c<c_{l}\approx 1.5$%, the conductivity plateau value depend on $c$ (Fig. 14(e)). This effect results from interlayer hopping, whereas for large $c$ ($c>c_{l}$) the conductivity plateau value is independent of $c$ like in MLG, and $\sigma_{M}=2\sigma_{M,{MLG}}$, with $\sigma_{M,{MLG}}=0.6$ G0 (Fig. 14(f)). As explained in [358], this transition from BLG to MLG behavior when the defect concentration $c$ increases, can be understood by comparing the elastic scattering length $L_{e}$ and the average traveling distance, $l_{1}\approx 2$ nm, in a layer between two interlayer hoppings.
    We also calculate the conductivity versus the inelastic length $L_{i}$. In the localization regime (very low temperature), as expected by the Anderson localization theory for 2D materials [189], we get:

### v2

@@ 3 Flat bands due to static defects
Electronic transport in graphene (see for example Refs. Berger04 , Novoselov2005 , Zhang05 , Berger06 , Castro09_RevModPhys , Yang18 ) and and other graphene-related 2D materials is sensitive to static defects that are, for example, frozen ripples, screened charged impurities, or local defects like vacancies or adsorbates. Adsorbates, which can be adatoms or admolecules (H, OH, CH3, …) attached to the surface of graphene, are of great interest in the context of functionalization to control the electronic properties Peres06 , Pereira06 , Pereira08a , Lherbier08 , Robinson08 , Wehling10 , Yuan10b , Bostwick09 , Leconte10 , Skrypnyk11 , Ferreira11 , Leconte11 , Lherbier11 , Lherbier12 , Trambly13 , Roche13 , Gargiulo14 , Xu16 , Jorgensen16 , VanTuan16 , Yang18 , Son20 . In particular, a functionalization by an adatom (or admolecule) covalently bound to an atom of the 2D material can be a resonant scatterer for conduction states which strongly affect electronic structure and transport properties.
@@
This appearance of a resonant state by functionalization, so-called midgap states, is particularly important in the case of bipartite lattices, i.e., systems in which atoms of a sub-network $\alpha$ are bound to atoms of a sublattice $\beta$, and reciprocally. Considering a first-neighbor tight-binding (TB) Hamiltonian, when the numbers of $\alpha$ orbitals and $\beta$ orbitals are different, isolated states appear at the energy $E=0$ (center of the band) Brouwer02 , Peres06 , Pereira06 , Pereira08a . Such a mechanism is exact globally for the whole structure. It also allows us to understand the appearance of resonant scatterers through a random functionalization, even if the number of $\alpha$ and $\beta$ orbitals is nearly the same. Indeed let us consider a random distribution of defects (adatoms or admolecule), with a not too strong concentration $c$ of defects. Locally, typically around a defect, an unbalance occurs between the two sublattices and a confined state appears at $E=0$ around the defect. When defects are in the same sublattice, these states are coupled to each other and form a kind of midgap band, leading to a peak in the DOS of width that depends on the concentration of defects. This mechanism applies to all the 2D materials presented in this review that have a bipartite lattice. Note that this yields also a simple picture to understand the appearance of edge states in a bipartite lattice, since an edge can produce a local unbalance between the two sublattices.
@@
Theoretical studies of transport in the presence of local defects have dealt mainly either with the Bloch-Boltzmann formalism (semi-classical methods) (see for instance Guinea08 , Castro09_RevModPhys , DasSarma11 , McCann13 ). In these theories, a major length scale that characterizes the electron scattering is the elastic mean free path $L_{e}$. These approaches indeed explain some experimental observations such as the quasi-linear variation of conductivity with a concentration of charge carriers. Yet, these theories have important limitations and can hardly describe in detail the localization phenomena that have been reported in some experiments. Indeed, in the presence of a short-range potential, such as that produced by local defects, the electronic states are localized on a length scale $\xi$. A sample will be insulating unless some source of scattering, like electron-phonon interaction, leads to a loss of the phase coherence on a length scale $L_{i}<\xi$. Note that electron-electron interactions may also play a dephasing role but this occurs usually at the lowest temperatures and we do not discuss it specifically here Lee85 , Polini20 . Therefore, in addition to the elastic mean free path $L_{e}$, the inelastic mean free path $L_{i}$ and the localization length $\xi$ also play a fundamental role for the conductivity of graphene with adsorbates. Moreover, the study of transport by semi-classical methods is valid when the Fermi energy $E_{F}$ is far enough from the Dirac energy. However, for $E_{F}$ close to the Dirac energy, the effects of resonant scatterers can strongly modify the electronic structure (such as gap opening Missaoui18 , Jouda21 …). These effects are important when the resonant scatterer concentration (defect concentration) is large with respect to the charge carrier concentration; indeed, each resonant scatterer creates one midgap state. Since these quantum effects, beyond the semi-classical behavior, are extremely dependent on the type of functionalization, systematic theoretical studies are needed to understand experimental results and stimulate new experimental studies.
@@ 3.1 Quantum transport calculation and calculated quantities
We used the real-space Kubo-Greenwood method, developed by D. Mayou, S. N. Khanna, S. Roche, and F. Triozon Mayou88 , Mayou95 , Roche97 , Roche99 , Triozon02 , which has already been used to study quantum transport in disordered graphene, chemically doped graphene and bilayers (see for instance Lherbier08 , Zhu10 , Lherbier11 , Leconte11 , Leconte11b , Lherbier12 , Roche12 , Roche13 , Trambly13 , Missaoui17 , Missaoui18 , Omid20 , Jouda21 , Chen23 , Miranda23 , Kaladzhyan23 , guerrero24_tBLG_disorder_Kubo , Guerrero25 ), functionalized carbon nanotubes Latil2004 , IshiiInelastic2010 , Jemai19 , Fan20 , and many other systems Fan14 , Fan20 such as quasicrystals Triozon02 , Trambly17 , 2D molecular semiconductors Fratini17 , Missaoui19 , and perovskites Lacroix20 . This numerical method connects the dc-conductivity $\sigma$, $\sigma(E,t)=e^{2}n(E)\mathcal{D}(E,t)$, to the DOS $n$ and the diffusivity $\mathcal{D}$, $\mathcal{D}(E,t)=\lim_{t\rightarrow\infty}{\Delta X^{2}(E,t)}/{2t}$ (see A.1.1), where the average square spreading $\Delta X^{2}$ is calculated at every energy $E$ and time $t$ by using a polynomial expansion method (A.3). This numerical approach has the advantage of using efficiently the method in real space. It takes into account all quantum effects due to a random distribution of static scatterers in a very large supercell containing more than $10^{7}$ orbitals. Most of our calculations are performed in a supercell of $\times1000$ or $\times1500$ cells of the initial cell of the studied crystal (graphene (2 atoms), Bernal bilayer (4 atoms), TMD (3 atoms), phosphorene (4 atoms / layer)), with periodic boundary conditions. Such a calculation may be performed by the recursion method (Lanczos algorithm) where the Hamiltonian in real space is written as a tridiagonal matrix Lanczos50 , Cullum1984 , Pettifor , Dagotto1994 of dimension $N_{r}$. Typically, we use $N_{r}=1000$$3000$ and we checked that results do not change significantly when $N_{r}$ increases. One can then directly evaluate the DOS from the Lanczos coefficients Haydock1972 , Pettifor , Gagliano1987 , Dagotto1994 , Schenk08 . This method is known as recursion method and has been used in some of our papers Trambly11 , Trambly13 , Trambly14 , Missaoui17 , Missaoui18 , Omid20 , but it leads to a convolution of the DOS by a Lorentzian function which a small width $\epsilon$. Thus $\epsilon$ is a kind of energy resolution of the calculation. Usually we use $\epsilon=1$ or $5$ meV. However, for systems with a gap, to avoid the tail expansion of the Lorentzian function in the gap, it is more efficient to diagonalize the tridiagonal Lanczos matrix of dimension $N_{r}\times N_{r}$ and to evaluate the DOS by Gaussian broadening of the spectrum Lacroix20 , Lacroix_these . Most of the time, we use a Gaussian standard deviation of $\epsilon=5$ meV. Note that for energies $E$ that are not close to the gap, the two methods give almost the same results.
@@
At finite temperature $T$, inelastic scattering caused by electron-phonon interactions are implemented by using the approximation of Relaxation Time Approximation (RTA). The details of the implementation of the RTA are given in Trambly13 (A.1.2). The conductivity $\sigma$ and the inelastic mean free path $L_{i}$ in the $x$-direction are thus calculated at every energy $E$ and every inelastic scattering time $\tau_{i}$. $\tau_{i}$ is the time beyond which the velocity autocorrelation function (Appendix Eq. (65)) goes exponentially to zero, and $L_{i}(E_{F},\tau_{i})$ is the typical distance of propagation during the time interval $\tau_{i}$ for electrons at energy $E$. Thus $L_{i}$ is also the distance beyond which a wave packet loses its phase coherence due to inelastic scattering processes, whereas elastic scattering events do not destroy the phase coherence. We know that $L_{i}$ decreases when the temperature $T$ increases, however the exact dependence of $L_{i}$ versus $T$ is unknown. This is why we consider different cases according to different possible values of $L_{i}$. Indeed, three different transport regimes may exist depending on the value of $L_{i}$ with respect to the elastic mean free path $L_{e}$, which is the average distance between two elastic scattering events. $L_{e}$ can be calculated approximately using the phenomenological formula (70) of the Appendix. The different transport regimes are:
@@
    Localization regime. When $L_{i}\gg L_{e}$, multiple scattering effects (such as weak or strong localization) strongly affect the transport and the conductivity is “macroscopic” in the sense that it is established over large sample sizes. This happens at sufficiently low temperature $T$, and then $\sigma$ decreases when $L_{i}$ increases (i.e., $T$ decreases). At each energy $E$, the localization length $\xi$ can be extracted by extrapolation of the curve $\sigma$ versus $L_{i}$, when $\sigma(L_{i}=\xi)=0$.
@@
    Diffusive regime. For smaller values of $L_{i}$, since $L_{i}\gtrsim L_{e}$, i.e., at higher temperatures (typically room temperature), $\sigma(L_{i})$ reaches a conductivity “plateau” close to the maximum $\sigma$ value (see A.1.2),
@@
    This regime is called “diffusive” regime, where $\sigma(L_{i})$ is almost independent of $L_{i}$ over a large $L_{i}$ range depending on the energy $E_{F}$. Usually the conductivity plateau corresponds to values of $L_{i}$ from a few nm to a few 10 nm, which may correspond to high temperatures and room temperatures, respectively. In this case, the conductivity of a sample depends only on quantum scattering over small distances which are typically of the order of magnitude of the distances between static defects ($L_{e}$); this is the reason why $\sigma_{M}$ is called “microscopic” conductivity. At each energy, the microscopic diffusivity $\mathcal{D}_{M}$ and microscopic conductivity $\sigma_{M}$ are defined as the maximum value of $\mathcal{D}(\tau_{i})$ and $\sigma(\tau_{i})$, respectively. The corresponding diffusivity is also called the microscopic diffusivity, $\mathcal{D}_{M}$, and it is possible to deduce from it the elastic mean free path $L_{e}$ (Eq. (70)).
@@
    Ballistic regime. The situation $L_{i}<L_{e}$ is an extreme case that most of the time should not be reached in real materials at realistic temperature (room temperature). This corresponds to the case of very pure materials with very few static defects. The conductivity is independent of static defects, and thus $\sigma(L_{i})$ increases when $L_{i}$ increases.
@@ 3.2 Mono and bilayer graphene with random distribution of defects
We start from the TB Hamiltonian,
@@
where $t_{ij}$ are the hopping terms between $p_{z}$ orbitals $i$ and $j$, and $\langle i,j\rangle$ is the double sum on index $i$ and $j$ with $i\neq j$. We can include only the first-neighbor hopping terms (first-neighbor TB: intralayer hopping $t_{ij}=-t_{0}=-2.7$ eV and for Bernal bilayer graphene (BLG) interlayer hopping $t_{ij}=t_{1}=0.34$ eV), or hopping terms beyond the first neighbors which is often more realistic using the same SK-TB parameters as in Sec. 2.3.2. Both models correctly recover the monolayer graphene (MLG) and Bernal bilayer graphene (BLG) energy band dispersions in the vicinity of the Dirac energy level $E_{D}=0$ Trambly12 , Missaoui18 . For MLG and BLG, the spectrum is symmetric with respect to $E_{D}$ (electron-hole symmetry) for first-neighbor TB (see Sec. 3.3.1), whereas it is slightly asymmetrical due to coupling beyond the first neighbors with the SK-TB (Sec. 2.3.2). It is important to note that another TB model exists for the BLG, in particular a well-known model deduced from the TB model for bulk graphite (see Ref. McCann13 and Refs therein). For the properties that we are studying here, the latter gives results comparable to the SK-TB model (Sec. 2.3.2).
@@ 3.2.1 Monolayer graphene (MLG)
Electronic properties in graphene with static defects such as vacancies, substitutions, or adsorbates has been intensively studied (see for instance Refs. Peres06 , Lherbier08 , Lherbier08b , Peres09 , Leconte10 , Lherbier12 , Roche12 , Roche13 , Trambly13 , Missaoui17 , Missaoui18 , Dutreix19 , Jouda21 ). We cannot go into all these studies in detail here, but we will just mention a few aspects that we consider essential. Roughly speaking, it is possible to propose a unified description of transport in one graphene sheet with adsorbates that fully takes into account localization effects and loss of electronic coherence due to inelastic processes Trambly13 . For this purpose, we distinguish between defects that are resonant scatterers or defects that are non-resonant scatterers. Examples of resonant scatterers are adsorbates that create a covalent bond with a carbon atom of a graphene layer, such as simple atoms or adsorbate molecules: H, OH, CH3… To simulate this covalent bond, we assume that the $p_{z}$ orbital of carbon, just below the adsorbate, is removed Pereira08a , Robinson08 , Wehling10 . As explained in the introduction of this part 3, in a monolayer graphene (MLG), which is a zero-gap material with a bipartite lattice, such functionalization creates so-called “midgap states” at the Dirac energy $E_{D}=0$. If we imagine a periodic structure with a low concentration of resonant defects, so that the defect states do not interact with each other, the midgap states would be associated with strict flat bands. These midgap states at energy $E_{D}$ are states that are only uncoupled from the rest of the band if we consider a Hamiltonian that includes only first-neighbor couplings. If couplings beyond the first neighbors are taken into account, the midgap states are coupled among themselves, and we then obtain a “midgap peak” whose width depends on the defect concentration, at an energy shifted with respect to $E_{D}$Trambly14 , Missaoui17 . This midgap peak is associated to flat bands in a periodic system. Examples of non-resonant adsorbates are adsorbates that create two covalent bonds with two neighboring C atoms of the MLG. This is the case for the adsorption of an oxygen atom (epoxy group). In generic terms, a non-resonant disorder can also be simulated by an Anderson disorder (random variation of the on-site energy of all orbitals). Such defects do not create resonant peaks, so they do not alter strongly the DOS, but they do affect the transport properties of charge carriers Trambly13 . As non-resonant defects are not associated with flat bands, we will not go into too much detail in this review.
@@ 3.2.2 Bernal bilayer graphene (BLG)
Figure 14: Electronic properties of Bernal AB bilayer graphene (BLG). (a) Sketch of the crystal structure of BLG. Atoms A1 and B1 are on the lower layer (layer 1); A2, B2 on the upper layer (layer 2). First-neighbor TB: $V^{0}_{pp\pi}=-t_{0}=-2.7$ eV and $V^{0}_{pp\sigma}=t_{1}=0.34$ eV are the intralayer and interlayer first-neighbor hopping terms, respectively. BLG is a bipartite lattice between sublattices $\alpha=\{{\rm A}_{1},{\rm B}_{2}\}$ and $\beta=\{{\rm A}_{2},{\rm B}_{1}\}$. Total DOS of (b) Monolayer (MLG) and (c) BLG with random distribution of vacancies with concentrations $c$ (%) of vacant atoms: (dashed line) First-neighbor TB (see panel (a)), and (solid line) SK-TB with hopping beyond first neighbor (see Sec. 2.3.2). The black dashed curve is pristine BLG ($c=0$), calculated with first-neighbor TB. (d) Conductivity $\sigma$ (solid line) and inelastic scattering length $L_{i}$ (dashed line) in BLG versus inelastic scattering time $\tau_{i}$ for concentration $c=1\%$ of resonant adsorbates (vacancies), and for 3 energy values, calculated with first-neighbor TB. (e,f) Microscopic conductivity $\sigma_{M}$ calculated with first-neighbor TB: (e) $c<c_{l}\sim 1.5$% and $\sigma_{M}$ with a constant minimum value (“plateau”) which decreases as $c$ increases, (f) $c>c_{l}\sim 1.5$% and $\sigma_{M}$ reaches a minimum value independent of the value of $c$ (see text). Concentration of vacancies: semi circle $c=0.05\%$, cross $c=0.25\%$, triangle up $c=0.5\%$, circle $c=1\%$, rectangle $c=2\%$, square $c=3\%$, triangle down $c=5\%$, star $c=10\%$. Insert shows enlarged curves around $E_{D}=0$. $G_{0}={2e^{2}}/{h}$. For the numerical method see A.3. From Missaoui17 , Missaoui_these .
@@
The unit cell of BLG contains four carbon atoms, A1, B1 in layer 1 and A2, B2 in layer 2 (Fig. 14(a)). Atoms A have three B first neighbors in the same layer and one A neighbor in the other layer, while atoms B have only three A first neighbors in the same layer. Since BLG is a bipartite lattice (see Sec. 16 below), an isolated functionalization by resonant scatterers creates a midgap state like in monolayer graphene (MLG). We first consider now the case of a random distribution of resonant defects in the two layers in BLG Missaoui17 . Figure 14 shows the total density of states (total DOS) $n(E)$ for different concentrations $c=0.5\%$ to $10\%$ of defects in MLG and BLG.
@@
Without defects, $c=0$, MLG and BLG DOSs differ Trambly13 , Missaoui17 for energies $E$ such as $t_{1}<E<t_{1}$ McCann13 , where $t_{1}$ is the interlayer hopping term (Fig. 14(b,c)). For small $c$ concentrations, $c<c_{l}\approx 1.5\%$, this distinction is still observed. Nevertheless, MLG and BLG have remarkably similar DOS for concentrations of defects $c$ larger than $c_{l}$. With the TB model (first-neighbor hopping only), midgap states occur at energy $E_{MLG}=0$. With the TB model including all neighbors (SK-TB), the midgap state is no longer at $E=0$, but it becomes a broad peak at negative energy $E_{MLG}$Missaoui17 . The value of $E_{MLG}$ varies from $E_{MLG}\approx-0.2\,{\rm eV}$ to $E_{MLG}\approx-0.3\,{\rm eV}$ when the defect concentration increases from $c=0.5\%$ to $10\%$, as in monolayer graphene Castro09_RevModPhys , Trambly14 . For large values of energies $E$ the DOS is weakly affected by the presence of disorder and near the Dirac energy there is an intermediate regime where the pseudogap is filled (“plateau”). We distinguish three cases according to energy values, which correspond to different transport regimes (Fig. 14) Missaoui17 :
@@
    For energies $E$ in the DOS “plateau” due to vacancies but not in the midgap states (Fig. 14(a)), the microscopic conductivity ($\sim$ room temperature conductivity) $\sigma_{M}$ (Fig. 14(e,f)) shows a “plateau” independent of $E$. For small $c$, $c<c_{l}\approx 1.5$%, the conductivity plateau value depends on $c$ (Fig. 14(e)). This effect results from interlayer hopping, whereas for large $c$ ($c>c_{l}$) the conductivity plateau value is independent of $c$ like in MLG, and $\sigma_{M}=2\sigma_{M,{MLG}}$, with $\sigma_{M,{MLG}}=0.6$ G0 (Fig. 14(f)). As explained in Missaoui17 , this transition from BLG to MLG behavior when the defect concentration $c$ increases, can be understood by comparing the elastic scattering length $L_{e}$ and the average traveling distance, $l_{1}\approx 2$ nm, in a layer between two interlayer hoppings.
    We also calculate the conductivity versus the inelastic length $L_{i}$. In the localization regime (very low temperature), as expected by the Anderson localization theory for 2D materials Lee85 , we get: